Can someone help me solve this problem?
The width of a swimming pool is 18.5 ft., and the length os 10.75 ft. I need to write the ratio of the length to width as ratio of whole numbers.
Thanks,
Lyn

Answer by fractalier(6550)

Length to width =
10.75 to 18.5 =
10.75 / 18.5
Let us multiply top and bottom by 4 to eliminate decimals...
4(10.75) / 4(18.5) =
43 / 74